OpenStudy (anonymous):

why is y²=x²-x⁴ <=> y= +/- x*sqrt(1-x²) ? what's the step I'm missing?

OpenStudy (anonymous):

what I don't see is how x comes out of the square root

OpenStudy (anonymous):

\[\sqrt{x^2-x^4}=\sqrt{x^2(1-x^2)}=\sqrt{x^2}\sqrt{1-x^2}=x\sqrt{1-x^2}\] assuming x is positive

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[1-x ^{2}\ge 0,1 \ge x ^{2} ,or x ^{2} \le 1,-1 \le x \le 1 ,y=\pm x \sqrt{1-x ^{2}}\]